OpenStudy (anonymous):

Find the slope of the straight line that passes through (–2, –4) and (3, –5).

OpenStudy (anonymous):

a.M=-1/5 b.M=-1 c.M=1/5 d.M=1

OpenStudy (unknownrandom):